### v2

# Evolutionary Optimization of Physics-Informed Neural Networks: Survey and Prospects

Abstract
Deep learning models trained on finite data lack a complete understanding of the physical world. On the other hand, physics-informed neural networks (PINNs) are infused with such knowledge through the incorporation of mathematically expressible laws of nature into their training loss function. By complying with physical laws, PINNs provide advantages over purely data-driven models in limited-data regimes. This feature has propelled them to the forefront of scientific machine learning, a domain characterized by scarce and costly data. However, the vision of accurate physics-informed learning comes with significant challenges. This review examines PINNs for the first time in terms of model optimization and generalization, shedding light on the need for new algorithmic advances to overcome issues pertaining to the training speed, precision, and generalizability of today’s PINN models. Of particular interest are the gradient-free methods of neuroevolution for optimizing the uniquely complex loss landscapes arising in PINN training. Methods synergizing gradient descent and neuroevolution for discovering bespoke neural architectures and balancing multiple conflicting terms in physics-informed learning objectives are positioned as important avenues for future research. Yet another exciting track is to cast neuroevolution as a meta-learner of generalizable PINN models.
Index Terms:
Physics-informed neural networks, physics-informed neuroevolution, evolutionary optimization, generalizability, survey.
I Introduction
Over the past decade, advances in deep learning have garnered increasing attention for their potential to address pressing challenges in science and engineering. The emerging topic of artificial intelligence (AI) for science promises to transform technological frontiers [1]. For example, learned models have shown the capacity to accelerate the design and discovery of novel materials, photonic devices, proteins, drugs, and renewable energy technologies such as wind turbines. However, unlike conventional applications of AI in vision, speech, or natural language processing where there is a deluge of available data, applications in scientific domains are often characterized by extreme data scarcity. This has raised concerns regarding the veracity of such models when trained only on small data. The cost of acquiring quality data gives rise to a practical bottleneck, leading to naive statistical models which sometimes generate physically inconsistent outputs.
The shortcomings of standard data-driven models have seeded the field of scientific machine learning where minimizing deviations from the known laws of nature serves to boost model precision and data efficiency [2, 3]. Associated methods offer unique pathways to achieving rationalizable AI, defined in [4], that is cognizant of fundamental physics principles. A myriad of approaches to instil physics consistency into learned models can be crafted, with the most suitable methods being naturally dependent on the task at hand and the type of knowledge available [5, 6, 7, 8]. Some instantiations include the design of model architectures that preserve invariances or symmetries representative of the underlying system [9], data augmentation with theory-based synthetic data (e.g., first principles numerical simulations) [10], or introduction of learning biases into the training process [11].
This review focuses on the growing body of literature on physics-informed neural networks (PINNs), popularized in recent years by the work of Raissi et al. [12]. The essential idea of a PINN—where mathematically-defined learning biases are directly incorporated into a neural network’s loss function—is simultaneously decades old, with early work in the 1990s [13, 14], and also a breath of fresh air for AI. This physics-informed loss is amenable to various forms of scientific knowledge and theories, including conservation laws and/or governing equations in the form of ordinary and partial differential equations (ODEs and PDEs), yet retains immense flexibility in the choice of network architectures (e.g., multilayer perceptron, convolutional or graph neural networks) and learning algorithms (e.g., gradient-based or gradient-free). This makes it possible for scientific knowledge uncovered over centuries, from the micro-scale evolution of quantum-mechanical systems to the macro-scale behavior of solids, fluids and propagation of electromagnetic fields, to be incorporated into the learned models of today.
Figure 1: Google Scholar survey of the literature shows the popularity of PINN over the years between 2019 and 2024, and PINN studies in diverse scientific areas. Data collected in November 2024.
TABLE I: Potential Applications of PINNs Across Various Scientific Domains
Branch Applications Reference
PINN for Fluid Mechanics Turbomachinery Design [15]
Airfoil Design [16]
PINN for Solid Mechanics Topology Optimization [17, 18]
Vehicle Crashworthiness Design [19]
Additive Manufacturing [20, 21]
Digital Twin [22, 23]
PINN for Chemistry & Electrochemistry Battery Management Systems [24]
PINN for Quantum & Particle Physics Diffuse Optical Tomography [25]
Nuclear Fusion Physics Analysis [26, 27]
PINN for Material Science Composite-Tool Systems during Manufacture [28]
In theory, PINNs compensate for a lack of labeled data by augmenting deep learning with physics principles. The potential of such physics-informed models has spawned significant interest across various scientific domains. Trends from a Google Scholar survey of the literature over 6 recent years (2019-2024) are depicted in Fig. 1. The figure reveals rapid growth in the popularity of PINNs in different branches of science, with many demonstrations being developed using open software libraries such as JAX [29], NVIDIA Modulus [30] (previously referred to as SimNet [31]), SciANN [32], and DeepXDE [33].
Despite the broad applicability of PINNs (ref. to Table I for a detailed summary of potential applications of PINNs), the switch from standard data-driven loss functions to physics-informed learning objectives has brought unforeseen difficulties in optimizing the latter’s uniquely complex landscape [34, 35, 36, 37, 38]. These training difficulties undermine the theoretical advantages of PINNs, and are currently a barrier to their adoption in place of established numerical methods in science and engineering. Existing surveys [8, 39, 40, 41, 42, 43] review PINN methods, applications, benchmark data, limitations, trends and prospects from interdisciplinary viewpoints, but a common point of emphasis is that efficient optimization of PINNs remain an open question direction [43].
Separately, despite the physics-informed loss, the physical validity of a PINN is limited to scenarios encountered during training, even after a complicated training process. They do not inherently generalize the learned physics to new scenarios beyond their training regime—e.g., changing PDE parameters, initial conditions or boundary conditions—thereby continuing to suffer from generalizability issues similar to standard data-driven learning algorithms.
In light of the above, this review paper examines PINNs for the first time through the lens of their optimization and generalization challenges. Notably, most PINN implementations today make use of gradient-based optimization algorithms. The ubiquity of the method may be attributed to inertia from the primacy of stochastic gradient descent (SGD) in conventional data-rich deep learning tasks. However, SGD is no silver bullet. Relative to standard deep learning, physics-informed loss functions are not only highly non-convex but also exhibit disparities in the magnitudes of back-propagated gradient signals between the data and physics loss terms (see II-A). These issues lead to pathologies that hinder effective training [44, 45]. The resulting optimization trajectory is littered with spurious minima, making it extremely challenging to find globally optimum solutions that satisfy physics. There is a high chance for gradient descent to get trapped in a local minimum that fails to satisfy the requisite physics, a possibility discussed in greater detail in Section III.
Figure 2: Boosting scientific ML with synergistic physics-informed neuroevolution (PINE) learning.
In order to fulfill the vision of data-efficient scientific machine learning, this paper asserts the need for algorithmic advances beyond gradient descent to move the needle for PINNs. Fig. 2 provides a schematic summary of some of the main computational challenges in this regard, and some corresponding proposed solutions. In Section IV, arguments are presented in support of our main claim: that algorithmic ideas developed within the computational intelligence community may be uniquely well-suited for physics-informed learning. Of particular interest are the gradient-free algorithms of neuroevolution that suffer less from spurious local minima due to their exploration-oriented global search capacity [46, 47]. In addition to optimizing network parameters, neuroevolution enables the discovery of bespoke architectures for physics-informed learning [48]. Evolutionary mechanisms could also prove useful for guiding training data generation in areas with higher errors on governing physics equations, having shown evidence of improved convergence to correct solutions [49].
An associated research direction is the synergy of neuroevolution with gradient-based training (in the spirit of memetic algorithms), marrying the best of both worlds. While population-based neuroevolution excels at navigating highly non-convex loss landscapes and balancing trade-offs between conflicting terms in physics-informed learning objectives, local search via gradient descent can facilitate rapid convergence to precise solutions. Yet another exciting track is to cast neuroevolution as a meta-learner of generalizable PINN models [50]. This may be achieved through a bilevel formulation where an upper-level evolutionary algorithm searches for PINN configurations that can be quickly fine-tuned for various downstream physics tasks at the lower-level.
The advancement of physics-informed neuroevolution is expected to have far-reaching impacts on establishing PINN methodologies at the forefront of scientific machine learning. Yet, surprisingly, there are only a handful of related works in the literature that explore such ideas, implying huge scope for future research and development.
II Overview of PINN Methodologies
PINNs are generally designed to learn a mapping between the input and output variables while satisfying specified mathematical constraints that represent the physical phenomenon or dynamical process of interest [12]. Given their ubiquity in modern science and engineering, PDEs are one of the most studied physics priors integrated in PINNs. Without loss of generality, we consider a physics-informed learning problem with input variables $x$ (spatial dimension) and $t$ (time dimension), and an output variable of interest $u$ satisfying PDEs of the general form:
PDE: $\mathcal{N}_{\vartheta}[u(x,t)]$ $=h(x,t),$ $x\in\Omega,t\in(0,T]$ (1a)
IC: $u(x,t=0)$ $=u_{0}(x),$ $x\in\Omega$ (1b)
BC: $\mathcal{B}[u(x,t)]$ $=g(x,t),$ $x\in\partial\Omega,t\in(0,T]$ (1c)
where the general differential operator $\mathcal{N}_{\vartheta}[u(x,t)]$ contains PDE parameters $\vartheta$ and can include linear and/or nonlinear combinations of temporal and spatial derivatives of $u$, and $h(x,t)$ is an arbitrary source term in the domain $x\in\Omega,t\in(0,T]$. Equation 1b specifies the initial condition (IC), $u_{0}(x)$, at time $t$=00, and Equation 1c specifies the boundary condition (BC) such that $\mathcal{B}[u(x,t)]$ equates to $g(x,t)$ at the domain boundary $\partial\Omega$, where $\mathcal{B}[\cdot]$ can either be an identity operator (Dirichlet BC), a differential operator (Neumann BC), or a mixed identity-differential operator (Robin BC).
Figure 3: PINN handles both scientific simulation and scientific discovery. Example applications are: 1) accurate simulation of transient vortex shedding behind a cylinder obtained by learning a PINN for a single problem solely using the governing physics (2D transient N-S equations); and 2) inference of the wave propagation speed, $c$, given sparse observations (locations demarcated by $\boldsymbol{\times}$), by learning an inference PINN model with a partially specified 1D transient wave equation.
II-A Physics-Informed Loss
PINNs are distinct from DNNs in that additional physics priors are incorporated as an essential component of its physics-informed loss function:
$\mathcal{L}_{PINN}=\lambda_{physics}\ \mathcal{L}_{physics}+\lambda_{data}\ \mathcal{L}_{data},$ (2)
where the relative weights $\lambda_{physics}$ and $\lambda_{data}$ control the trade-off between the two components. The physics-informed component can consist of multiple PDE, IC, and BC loss terms:
$\mathcal{L}_{physics}$ $=\lambda_{pde}\ \mathcal{L}_{pde}+\lambda_{ic}\ \mathcal{L}_{ic}+\lambda_{bc}\ \mathcal{L}_{bc}$ (3a)
$\mathcal{L}_{pde}$ $=\lVert\mathcal{N}_{\vartheta}[f(\cdot;\boldsymbol{w})]-h(\cdot)\rVert_{L^{2}(\Omega\times(0,T])}^{2}$ (3b)
$\mathcal{L}_{ic}$ $=\lVert f(\cdot,t=0;\boldsymbol{w})-u_{0}(\cdot)\rVert_{L^{2}(\Omega)}^{2}$ (3c)
$\mathcal{L}_{bc}$ $=\lVert\mathcal{B}[f(\cdot;\boldsymbol{w})]-g(\cdot)\rVert_{L^{2}(\partial\Omega\times(0,T])}^{2}$ (3d)
where $f(\cdot;\boldsymbol{w})$ is the PINN output function with network parameters $\boldsymbol{w}$, evaluated on a set of label-free training points (collocation points) sampled from the respective spatio-temporal domain. An illustrative schematic of a PINN model is presented in Fig. 3. Evaluation of differential operators, e.g. $\frac{d\ f(\cdot;\boldsymbol{w})}{dt}$, $\frac{d\ f(\cdot;\boldsymbol{w})}{dx}$ and $\frac{d^{2}f(\cdot;\boldsymbol{w})}{dx^{2}}$, are generally required for computation of the physics-informed loss on these training samples. They can be conveniently computed by automatic differentiation [12] and/or numerical differentiation [51, 52] methods at any sample location.
Fundamentally, PINNs arrive at an accurate and physics-compliant prediction for a desired physics scenario by forcing its outputs $f(\cdot;\boldsymbol{w})$ to satisfy Equation 1 through minimizing the physics-informed loss (Equation 3). The effectiveness of the optimization of network parameters $\boldsymbol{w}$ through the physics-informed loss is a primary determinant of the accuracy of the PINN model. In contrast, the typical data-driven loss (as common in a DNN) computes the MSE between the model prediction $f(x_{i},t_{i};\boldsymbol{w})$ and available target $u_{i}^{label}$ over $i$=$1,...,n$ labeled samples:
$\mathcal{L}_{\text{DNN}}=\mathcal{L}_{data}=\frac{1}{n}\sum_{i=1}^{n}\left(u_{i}^{label}-f(x_{i},t_{i};\boldsymbol{w})\right)^{2}$ (4)
Figure 4: Illustrative example showing how a PINN model can make accurate prediction by learning with physics laws (1D convection-diffusion equation) and few labeled data, whereas a conventional DNN model fails to learn the correct output.
II-B Physics-Informed Learning Advantages
The advantages of physics-informed learning are demonstrated in Fig. 4. Given only a few labeled data ($n$=$5$), a DNN model (with sufficient representation capacity) remains unable to correctly reconstruct the entire system, although it fits very well on all training samples. On the other hand, a PINN model can accurately learn the solution from the same set of labeled data. Given that the physics constraints are enforced on a set of label-free training samples ($n_{pde}$=$1000$) uniformly drawn from the computation domain utilizing the aforementioned PINN loss, a PINN can learn the solution even without any labeled data. The governing physics in this example is the convection-diffusion equation $\alpha\frac{du}{dx}-\frac{d^{2}u}{dx^{2}}=0$ with PDE parameter $\alpha$, which is a ubiquitous physics model in science and engineering that describes the distribution of a scalar quantity (e.g., mass, energy, temperature) in the presence of convective transport and diffusion [53].
II-C PINN for Scientific Simulation
PDEs offer concise mathematical descriptions of diverse physics phenomena in nature. With PINNs, it is possible to simulate an entire physical system by training the model to approximate the solution of the PDEs specific to the target physics, without any labeled data (i.e., solely using the physics-informed loss). A good example of this is the fluid dynamic simulation of vortex shedding behind a cylinder depicted in Fig. 3, which was obtained by using a PINN to solve the transient 2D Navier-Stokes (N-S) equations [54].
Compared to classical numerical solvers employing the finite difference, finite element, or finite volume method [55], the PINN approach has the significant advantage of being mesh-free. Meshing, which is the process of domain discretization, is an essential, yet non-trivial task. Classical numerical methods can fail to find the correct solution if the meshing is not appropriately done. On the other hand, PINNs can provide a solution without similar mesh requirements. Additionally, the solution obtained through PINN is differentiable, and its compact representation requires lower memory demand for storage. This makes PINNs a promising alternative to traditional numerical methods for simulating physics phenomena; an extensive review is presented in [39].
While much of the PINN literature focuses on single problems, a PINN model can also incorporate additional parameterizations in the input space, allowing for learning from and solution across a parametric family of physics scenarios [56, 57, 58, 59]. These parameterizations can include geometrical aspects or physically meaningful system parameters, such as the Reynolds number or material dielectric constant. These models, also referred to as parameterized PINN, can then provide predictions across a set of physics scenarios as opposed to providing the solution for a single scenario.
II-D PINN for Scientific Discovery
In addition to directly simulating physical behavior, PINNs can be seamlessly extended to various kinds of hypothesis inference problems useful for scientific discovery. These problems typically involve an incomplete and potentially noisy set of observed data points, and the goal is to infer information about the system through optimizing the hypothesis space. The loss function used to train PINN models for inference consists of both physics-informed (to capture relevant physics) and data (observations) components (Equation 2). The hypothesis spaces of the inference problem are included as optimization variables along with the network parameters during training; see example in Fig. 3.
The PINN models can be applied to:
  1. 1.
    Inference of system-level parameters such as the underlying governing equations or constitutive relationships or system parameters (e.g., Reynolds number, wave speed) based on observations [60, 61, 62, 63, 64, 65].
  2. 2.
    Inference of unobserved field variables across the entire domain based on limited observations, e.g., predicting pressure in a particle imaging velocimetry experiment with only velocity observations [66]. Many have employed PINNs towards the inference of domain field variables in various problems spanning fluid dynamics [67], heat transfer [68] and biomedical sciences [69, 70].
  3. 3.
    PDE-constrained design optimization towards a set of desired (observation) outcomes. The PINN is used as an adjoint optimizer, enabling the design parameters to be optimized towards the desired outcome while fulfilling all specified relevant physics during training. This has been successfully demonstrated on topology optimization [71], metamaterial [72] and biomimetics [73] design problems.
PINN models are uniquely suited for hypothesis inference when compared to numerical solvers and can effectively synergize theory and observation data from various sources such as sensors and imaging, even for ill-posed PDE problems. It can be combined with probabilistic inference methods (e.g. Bayesian-PINN [74]), with the probabilistic framework offering a mathematical tool to handle uncertainties arising from both noisy data and physics assumptions [74, 75, 76].
Figure 5: The plots contrast the local loss landscapes of PINN and DNN, for the convection-diffusion problem presented in Subsection II-B. After 100,000 training iterations, we project their corresponding loss values onto the plane spanned by the first two principal Hessian directions as per previously reported literature [34] to get an intuition for the loss landscape.
III Optimization and Generalization Challenges
However, PINN’s advantages come at the cost of increased difficulty in optimizing PINN model parameters, with many early papers showing long training times and limited predictive performance on more complex problems [12, 36, 37, 38, 34, 35]. Moreover, parameterized PINNs can still face issues with generalizability since the accuracy of interpolation-based predictions on a new problem is not guaranteed, and training a well-generalizable PINN model may not be an easy task. In this section, we shed light on some contributing factors to the optimization and generalization challenges—2 pre-dominant hurdles to the wide-spread adoption of PINNs in real-world applications today—and how they may originate from fundamental limitations in gradient-based approaches.
III-A Challenges in Optimizing PINN Model
In many studies involving PINN, it was observed that the training iterations required for convergence far exceed those needed for typical data-driven DNN. Unsuccessful PINN training (premature convergence) can lead to unacceptable (unphysical) simulation outcomes. We relate the optimization difficulty of PINNs to a few factors below.
III-A1 Complex Loss Landscape
The physics-informed loss usually consists of multiple PDE and IC/BC loss terms (Equation 3). The different loss terms can display distinct scales and characteristics during PINN training. For instance, the scale of PDE loss can span several orders of magnitude depending on the underlying differential operators and the complexity of the solution, and this range is only revealed during training. Additionally, some loss terms change more dramatically than others during training, as shown by prior analyses through the lens of neural tangent kernel (NTK) [77, 78, 79]. These characteristics make the proper scaling of different loss terms challenging.
In addition, the PINN optimization landscapes are littered with spurious minima and are more complex than those of usual data-driven DNN models. Fig. 5 illustrates the loss landscapes of PINN and DNN models for the convection-diffusion problem presented in Subsection II-B. This visualization provides evidence of a much higher degree of complexity and ruggedness of the loss landscape for PINN compared to usual data-driven DNN models. The smoothness of the latter is conducive to gradient-based algorithms that can easily find descent paths from the point of initialization to a good local optimum that fits the training data well. This outcome corroborates the primacy of SGD and its variants in today’s deep learning tasks. In contrast, the contradictory descent directions from different loss terms encountered during PINN training can adversely impact optimization [80, 81].
Although these powerful gradient-based algorithms are inseparable from the recent success of deep learning, they are no silver bullet. Gradient-based methods are susceptible to being trapped in spurious local minima, which is often the case in the context of PINN optimization. Moreover, usual deep learning applications do not require a fully converged solution because they prioritize generalization performance over training precision. However, this is not the case for PINN, where the loss needs to be as close to the global minimum as possible to ensure the veracity of the model concerning the underlying physics of a problem. A small violation of the PDE constraints can result in an unphysical PINN solution while a small violation of the BC or IC constraints corresponds to an entirely different PDE problem [82, 83]. This issue is especially detrimental in time-varying dynamical problems where small violations early in time can cause errors to propagate and accumulate in later times.
III-A2 Initial Local Minimum Trap and Spectral Bias
PINN models using common weights initialization (such as Xavier initialization) and activation function (such as tanh, sigmoid) are initialized around flat output functions, and are consequently susceptible to a local minimum trap at initialization whereby the PINN loss already sees a minimized PDE loss term; see theoretical analysis in [84, 85]. Such a strong bias around initial flat functions can be difficult for gradient descent optimization to escape, especially when sensible solutions that simultaneously minimize all the PDE and IC/BC loss terms are not accessible near the point of initialization.
Another closely related issue stems from the known saturation property of the activation functions commonly used for PINN, such as tanh and sigmoid. Whenever the network weights are shifted towards a larger distribution, e.g., as driven by the need to approximate higher frequencies and steep gradients common in physics, such saturation property can cause their backpropagation gradients to vanish. This makes further optimization of the network weights hard for gradient descent, causing a learning bias towards smooth functions. This known tendency—also called spectral bias—has been studied through the lens of NTK [86, 87], revealing a significant discrepancy in convergence rates for different frequency components in the output during training (lower frequencies are learned before higher frequencies). While this has been regarded as a natural regularizer with potential benefits to the generalization performance of DNNs, the consequences are highly deleterious for PINNs. Spectral bias makes it difficult for PINNs to accurately learn complex nonlinear physics behaviors [88], e.g., in the vicinity of steep gradients characterizing shock waves. These local discrepancies further deter propagation of the correct initial and boundary conditions into the computation domain, which has been suggested to be critical for successful PINN training [83, 89], and eventually becomes a bottleneck to convergence of the PINN loss.
This saturation also translates to increased difficulties for gradient-based algorithms to alter the behavior of PINN output far away from the input origin. As a result, PINN struggles to learn over long time horizons, partly because the learning bias towards smooth functions can cause the gradient descent to prematurely converge to a trivial solution (satisfying the PDE) before the initial and boundary conditions can properly propagate to a later time [90].
III-A3 Sensitivity to Sample Size and Location
The training samples of the PINN model are typically uniformly drawn from the input domain and are equally important everywhere due to their physical correspondence with the spatio-temporal extent of the problem. While the use of mini-batch samples is a standard procedure in gradient-based training, it remains an open question how one can efficiently reduce the batch size (and hence the training cost) while maintaining the desired PINN model accuracy [91]. Using a large batch size is crucial for the accuracy of PINNs due to the potential need to learn high frequencies or complex structures in the a priori unknown target function. However, this inevitably increases training costs. Furthermore, oversampling on a high loss region may adversely slow convergence since it can be harder to learn such local complex gradient behaviour—a cause of stiffness during training [51, 52, 37, 38]. On the other hand, an overly small batch size will result in noisy (big jump in loss values) and inconsistent local gradient signals [92], thereby causing issues during gradient descent.
III-B Challenges in Learning Generalizable PINN Model
With the simulation (hypothesis evaluation) and inference capability, PINNs are versatile and can be applied to various applications common in science such as model-predictive-control, scenario planning, and design exploration [93, 94, 95, 96]. The commonality to these applications is the need for multiple accurate and physics-compliant predictions across a series of related problems. These repeated evaluations usually differ by modifications to the physics scenario such as changes in geometric shape, PDE parameter, IC and BC.
III-B1 How Generalizable Is A PINN Model?
Most parameterized PINN models (ref. to Subsection II-C; they are designed to learn from a parametric family of physics scenarios) make predictions on new scenarios unseen during training through interpolation. Contrary to its promise, the interpolation is usually physics-agnostic, with similar negative implications for generalizability as the predictions from typical DNNs [50]. Given the difficulties in optimizing a PINN model for a single physics scenario, the cost of performing physics-informed learning across a set of scenarios can be prohibitive [97]. The effectiveness of these parameterized PINNs for prediction of new scenarios then depends on the complexity of the problem. Moreover, the problem’s parameterization needs to be determined at the learning stage, which reduces the flexibility of the parameterized PINN model when applied to new physics scenarios.
III-B2 How to Train A Generalizable PINN Model in The Data-Scarce Regime?
Meta-learning approaches provide a systematic framework for discovering good weight initializations that can be quickly adapted—with minimal training—to arrive at physically precise predictions in every new scenario (task) for a given problem distribution. Meta-learning of a generic PINN model initialization can also allow generalization across diverse, multi-modal problem distributions, hence it holds immense promise for addressing generalization challenge of PINN [98, 99, 100, 101, 97, 102, 103, 104, 105].
There are still a number of unresolved issues in achieving a generalizable PINN model via meta-learning. A fundamental assumption of popular gradient-based meta-learning algorithms—such as model-agnostic meta-learning (MAML) algorithm [106] and its variant, Reptile [107]—is that the models can learn well on each task-specific training after a few gradient descent updates from the initialization. However, this is usually not the case for PINNs because physics-informed losses are much more challenging to optimize. They require more extensive training to achieve good convergence [97]. With only a handful of task-specific gradient descent updates, gradient-based meta-learning algorithms may still be prone to getting stuck in local minima as the gradients collected from a batch of diverse training tasks can be noisy or even deceptive. As a result, recent PINN studies have shown inferior results for MAML, potentially due to insufficient task-specific adaptation and non-convergence in the meta-learning algorithm’s outer loop optimization [102, 103].
IV Physics-Informed Learning with Neuroevolutionary Algorithms
Despite their theoretical advantage in diverse scientific domains, Section III highlighted several practical challenges in today’s PINN learning and optimization. In this section, we postulate that many algorithmic ideas developed in the computational intelligence community may be uniquely well suited for addressing issues plaguing PINNs. One exciting research direction is to synergise physics-informed neuroevolution with local gradient descent or least-squares signals, thereby leveraging the best of both worlds to quickly arrive at physically accurate and generalizable PINN models. In what follows, we delve deeper into this and other algorithmic pathways deemed worthy of exploration.
IV-A Exploration-Oriented Evolutionary Algorithms with Global Search Capacity
IV-A1 Neuroevolution
While examining PINN optimization challenge, it becomes apparent that there is a need for new algorithms that can overcome issues such as local minima traps and vanishing gradients faced by conventional gradient descent algorithms, including the widely used SGD. Gradient descent is a point-based optimization method that is highly sensitive to the issues of local minima or learning bias magnified by PINN loss functions. In contrast, in silico evolutionary computation is a class of highly parallelizable approaches for guiding populations of candidate solutions towards near-optimal regions of search / parameter spaces via mechanisms inspired by biological evolution [108]. We shall broadly refer to any such algorithmic realization as an evolutionary algorithm (EA) and its application to the training of neural networks as neuroevolution.
The operators of EAs are typically based on principles of natural selection. They do not require derivative computations of the loss functions being optimized and are therefore gradient-free. Importantly, as population-based methods, EAs offer global search capability with lesser tendency to get trapped in a single local minimum [109, 47]. What’s more, since their search is typically invariant to order-preserving transformations of the loss function [110], EAs are less sensitive to small (vanishing) gradients. For these reasons, they are deemed as potential alternatives to SGD, particularly in the context of PINN where the goal is to converge to physically precise solutions with close to zero loss [46]. In addition to its global search capacity, an evolved population naturally provides an ensemble of neural networks that affords the possibility of robust predictions with quantified uncertainty [111, 112].
In neuroevolution, EAs are often tasked with directly optimizing the parameters of neural networks. However, the large-scale nature of neural network parameters poses significant challenges, as EAs typically struggle to perform effectively under such conditions. To address this, several advanced techniques, such as sparse optimization [113] and coevolution [114], have been developed to enhance the performance of EAs for large-scale variable optimization. They provide valuable insights and inspiration for addressing large-scale optimization challenges within the context of PINE.
IV-A2 Pareto Optimization
The existence of multiple loss terms—see Eqs. (2) and (3)—makes PINN training intrinsically a multi-objective optimization problem. Since the terms $\mathcal{L}_{data},\mathcal{L}_{pde},\mathcal{L}_{ic}$ and $\mathcal{L}_{bc}$ represent different physical quantities, their magnitudes may be incommensurable. In this case, the role of the scalarizing weights is not only to express preferences between the losses, but also to compensate for differences in their magnitudes. The possible conflict between the losses and the non-convexity of the aggregated physics-informed learning objective implies that a neural network, despite being a universal function approximator, is not guaranteed to find a solution that simultaneously minimizes all terms [44]. Moreover, the shape of the Pareto front formed by the multiple objectives is not known beforehand [115], raising concerns that a simple weighted-sum aggregation approach may not support all parts of the front [116].
Multi-objective EAs (MOEAs) are tailor-made to address the issues mentioned above [117]. In particular, Pareto dominance-based variants of MOEAs [118] do not require the multiple loss terms to be scalarized, and hence are less sensitive to incommensurable objectives. Their population-based search mechanisms also allow for free-form convergence towards the entire Pareto front in a single run of the algorithm, irrespective of the front’s shape. These salient features have naturally begun to attract the attention of PINN researchers, although associated developments are still at a nascent stage. In [119], the non-dominated sorting genetic algorithm [120] was hybridized with a variant of SGD (namely, Adam [121]) for training the parameters of a PINN. The exploration capacity of the EA was shown to help overcome local minima and enable precise satisfaction of the initial and boundary conditions encoded in the physics-informed loss. Given the computational cost of a hybrid framework, a pure population-based multi-objective search algorithm was studied for PINN training in [112], leading to calibrated uncertainty quantification by examining the empirical distribution of the ensemble’s predictions.
In addition to using MOEAs for PINN parameter training, there have been efforts towards evolving the model’s hyperparameters (including the scalarizing weights and activation functions). The method yields a Pareto set of hyperparameters where each setting offers an optimal trade-off between the data and physics losses. Practitioners can then decide post-hoc which setting they wish to pursue [122]. An analysis of such post-hoc decision-making revealed that solutions close to the convex bulge of the Pareto front typically led to performant PINNs [123]. Evolutionary algorithmists will immediately notice a connection between this observation and the literature on knee-point driven MOEAs [124], which could serve as a valuable resource for hyperparameter tuning in PINNs.
IV-A3 Memetic Algorithms
There is a clear theoretical rationale supporting physics-informed neuroevolution, yet there are only a handful of related works in the literature that explore such ideas. A barrier to the widespread use of EAs has been their high sample complexity, a consequence of updating and evaluating populations of solutions over multiple iterations. It is useful to consider hybridizing global evolutionary search with local gradient signals—thus inheriting the merits of both approaches. While the former can serve as an efficient basin-hopper for exploring rugged PINN landscapes, the latter facilitates rapid convergence to stationary points by exploiting local derivative information of loss functions.
There exists a long history of research in evolutionary computation studying synergies of this kind, with adaptive memetic algorithms forming a class of methods tailored for this purpose [125]. In the deep learning literature, gradient-guided evolutionary approaches have been designed for training neural networks [126, 127], with Cui et al. [128] proving a simple theorem that the performance of an evolutionary stochastic gradient descent method with elitist selection strategy never degrades with increasing iterations. However, to the best of our knowledge, there is no comparable body of work for PINN. Davi and Braga-Neto [111] were able to show that combining a swarm optimization algorithm (a population-based method like EAs) with gradient descent could produce an ensemble of PINN models that out-performed pure gradient-based learning. This result illuminates unexplored avenues for unifying SGD with gradient-free EAs as complementary algorithms within a single framework for finding globally optimum PINNs.
The crafting of synergistic optimization frameworks to meld gradient-based and gradient-free learning poses fundamental challenges, including determining interaction mechanisms that best harness the strengths of the two approaches. Key questions to address include balancing the intensity of SGD versus neuroevolution, ascertaining when gradient descent should take over from evolution, or selecting an appropriate subset of the evolving population for fine-tuning with gradient descent (as doing so for the entire population would be too expensive). Notably, the subfield of memetic computation in computational intelligence—where ‘memes’ are loosely defined as computationally-encoded units of information (e.g., local gradient signals) combined with global evolutionary search for effective problem-solving—offers a wealth of theories and methodologies that directly tackle these issues [129, 130].
IV-A4 Hardware-Accelerated Neuroevolution
Together with algorithm-centric research tracks, hardware-accelerated neuroevolution offers a promising path for EAs to substitute or be hybridized with SGD for PINN training. Sung et al. [131] provide preliminary results in this regard, demonstrating that neuroevolution with JAX [132] could lead to convergence trends that are orders of magnitude faster than standard implementations. The authors have made available codes for a set of five benchmark problems [^1] Python source codes for neuro-evolving PINN using EvoJAX are available at https://github.com/nicholassung97/Neuroevolution-of-PINNs., representative of real-world phenomena in classical mechanics, heat and mass transfer, fluid dynamics and wave propagation, with associated performance metrics and baseline results to support future research into the development of customized EAs for PINN training.
IV-B Neural Architecture and Hyperparameter Search
The choice of architectures and hyperparameters for PINNs has often been guided by experiential priors drawn from standard deep learning. However, a growing body of literature shows that accurate physics-informed learning may call for bespoke architectures. For example, contrary to popular belief, it was found that more hidden layers do not necessarily mean better performance and can sometimes be detrimental. An automated neural architecture search via a bilevel programming formulation revealed that, for certain governing differential equations, a shallow neural network with many neurons was more suitable for PINNs [133]. Similarly intriguing insights regarding the choice of activation function have also been uncovered. While squashing functions (e.g., tanh and sigmoid) and variants of the rectified linear unit are the mainstay in the deep learning literature, the use of Fourier features (i.e., sine and cosine) has gained prominence in physics tasks [84, 85], specifically for their ability to overcome spectral bias and capture periodic or high frequency patterns (e.g., shock fronts) common in various physics applications [134, 135, 136].
Given the large number of design choices for network architectures and hyperparameters, manual experimentation to identify an optimal design for every physical problem is impractical. Notably, there exists a rich history of research in neuroevolution that automates such design choices through the joint evolution of neural network topologies along with their parameters [137, 138, 139]. One such example of the fully automatic design of PINNs by a genetic algorithm with specially tailored selection, crossover, and mutation operators was proposed in [48]. Each chromosome of this method encodes a learning algorithm, architecture of the PINN model, activation function in the hidden layers, and the output activation. In a related approach, an asynchronous genetic algorithm that alternates between model selection and optimization of physics-informed neural state space dynamics models was presented in [140]. Daw et al. [89] proposed an evolutionary sampling procedure to incrementally accumulate sample points in regions of high PDE residuals for mitigating training failures.
In order to achieve higher approximation accuracy and faster convergence rates, an EA for concurrently exploring optimal shortcut connections between PINN layers and novel parametric activation functions expressed by binary trees was put forward in [141]. Interestingly, numerical studies show that random projection PINNs—where most network parameters are randomly initialized and fixed, with training limited to the output layer—significantly accelerate training on physics-informed learning objectives [142]. However, ensuring predictive accuracy in this setting requires careful configuration of the randomization bandwidth: a hyperparameter defining the width of the probability distribution from which the random parameters are sampled. In [143], this hyperparameter optimization was carried out by the differential evolution algorithm [144], resulting in PINNs that outperformed higher-order finite element methods (an established numerical method for solving differential equations) in terms of speed and accuracy.
Training a PINN model can be computationally intensive, rendering the evaluation of hyperparameters or network architectures prohibitively expensive. To address this problem, various surrogate-assisted evolutionary techniques [145] have been proposed within the field of evolutionary computing, emerging as promising solutions. These techniques encompass a wide range of optimization scenarios, including single-objective [146], multi-objective [147], large-scale [148], constrained [149], discrete variable  [150], and even mixed-variable optimization [151]. These techniques provide a strong foundation for developing more efficient hyperparameter optimization and architecture search methods tailored to PINNs.
Figure 6: The projectiles under the effect of Earth, Mars, and Moon are respectively solved by transfer neuroevolution—dubbed tNES. The convergence trends from SGD, xNES, and tNES optimization runs for selected scenarios are displayed (the bold lines indicate the mean convergence path of multiple independent runs and the shaded areas indicate their 10th-90th interpercentile ranges). Individual solutions given by multiple SGD and tNES (transfer from a relevant prior, i.e., source distribution obtained by solving a different gravity $g$ or launch angle $a_{0}$) optimization runs for one particular scenario ($g$=1.6, $a_{0}$=45) are plotted and compared against the ground truth.
IV-C Transfer Learning and Transfer Evolutionary Optimization
IV-C1 Learning PINN on Related Tasks
In a scientific study, it is common for a single set of PDEs to be evaluated under different physics scenarios (tasks), e.g., variations in PDE parameters, and IC/BCs. Moreover, for dynamical problems with causality, PINNs can generally be formulated to predict sequences of physical states, such as at different crack steps in a crack propagation study [152] or at different time steps in a time-varying dynamical problem [153, 154]. Such PINN models are sequentially learned using solutions from previous steps as IC/BCs. Relevant task solving experiences (i.e., optimizing PINN model parameters) can naturally accumulate over time. This allows transfer learning and optimization algorithms to exploit reusable knowledge from past episodes and jump-start the physics-informed learning, hence achieving rapid convergence for new target task [155]. The transfer learning of PINN from simpler tasks has also been shown to improve solution accuracy on challenging tasks, such as high-frequency and multi-scale problems, compared to learning the model from scratch [156].
Transfer learning in the context of PINN commonly involves transferring network weights from a similar physics scenario that has been previously learned (source task). These transferred weights serve as initialization for training the target task. Typically, only the last few layers of the networks are optimized after the transfer, allowing the sharing of previously learned useful features between target and source tasks. A singular value decomposition-based transfer approach has also been proposed in [157], which involves freezing the singular vectors of the dense kernel matrix of the hidden layers and only optimizing the singular values for the target task. Apart from network weights, the study from [158] has shown that transfer learning can also facilitate the learning of loss-balancing weights in both forward and inverse PINN models.
Another interesting approach worth exploring is to combine physics-informed learning and data-driven learning techniques using transfer learning. A multi-fidelity transfer learning framework was proposed in [159] to first train a PINN model to capture important information about the governing physics of the problem from the approximated (low-fidelity) physics equations. The model then undergoes data-driven learning with available high-fidelity data, leading to accurate predictions even with limited high-fidelity training data. While the study has demonstrated that transfer learning from physics can benefit a data-driven DNN, the reverse pathway deserves further investigation. There is potential for improvements to PINN optimization by harvesting the experience gained from a data-driven DNN model through transfer learning and transfer evolutionary optimization.
IV-C2 Adaptive Transfer
An important question in the transfer of physics-informed learning is the identification of relevant source task(s). If the source task is far from the target task, there could be insufficient commonalities between these tasks for the transfer to be beneficial, and it could even negatively impact the learning performance of the target task (negative transfer). For problems with causality, a natural choice for source is the networks learned on the previous time-step or physical state [153, 154]. However, the target function is generally a priori unknown in physics-informed learning, hence calculating the similarity between target and source tasks based on their output functions is not feasible. Identifying the most relevant source task then becomes a challenge that requires prior knowledge of how the system behaves when its geometry, PDE, and IC/BC parameters change [160]. This becomes even more challenging when dealing with previously solved tasks involving multiple parameter changes.
Adaptive transfer learning can be a particular strategy to address the issues without the need for explicit a priori identification of relevant sources. For example, an adaptive transfer physics-informed learning framework was proposed in  [160] to achieve effective transfer, even when there is significant variation in the PDE parameters between the source and target tasks. The approach treats the PDE parameters as trainable, hence, they can be adaptively updated from the source to the target parameter values, creating a low-loss path in the parameter space to guide the optimization from the source to the target task. Throughout this process, the loss is consistently kept at a low level (similar to the concept of a minimum energy path connecting two minima [161]), ensuring that the PINN solution remains very close to the physical solution and thus guaranteeing the stability of the training process.
A recent study by Wong et al. [46] on the transfer neuroevolution of PINN demonstrated that accurate solutions for new and harder PDEs can be rapidly achieved by optimizing PINN with EA. A key aspect of this transfer optimization approach is its ability to incorporate multiple sources of knowledge while optimizing the PINN for the specific target task. This can be achieved through a principled probability mixture model-based EA that adaptively assigns weights to the different sources based on their observed relevance to the target task during training, thereby preventing detrimental negative transfers. Fig. 6 illustrates the accelerated convergence trends achievable through the transfer neuroevolution of PINN, with clear optimization advantages relative to xNES (a state-of-the-art of EA [162]) and SGD-based (Adam) training.
Similar probabilistic formulations are also possible in multitask neuroevolution [163, 164], enabling the simultaneous solving of sets of related PDEs in a single algorithmic pass. A multitask approach could prove much faster than solving PDEs independently, as useful information is shared between tasks without the need to optimize from scratch.
Figure 7: Meta-learning PINN with MAML (left) versus Baldwinian evolution (right). In MAML, the initial PINN weights ${\bf\it W}$ are learned using gradient-based method, such that they can be quickly fine-tuned (physics-informed learning) on new tasks. As the task-specific physics-informed learning is limited to one or a few gradient descent updates during training, such limited fine-tuning is usually insufficient for a meta-learned PINN to converge. In Baldwinian evolution, the weight distribution in the pre-final nonlinear hidden layers ${\bf\it\tilde{w}}$ and learning hyperparameters $\lambda$ are jointly evolved. The task-specific physics-informed learning is performed on the output layer weights $w$ (segregated from ${\bf\it\tilde{w}}$) with a 1-step pseudoinverse operation at both training and test time (for linear ODE / PDEs).
IV-D Meta-Learning by Neuroevolution
IV-D1 Meta-Learning in PINN
Meta-learning algorithms provide a systematic framework for reusing knowledge from a distribution of learning tasks. This allows a meta-learned PINN model to quickly adapt to new physics-informed learning tasks with minimal training. Crucially, meta-learning offers a pathway towards a generalizable PINN over a broad class of physics scenarios, including different PDE families and geometries, as highlighted in Subsection III-B2. For example, the MAML and Reptile algorithms have been recently explored in the context of PINN [102, 103, 101]. They seek to meta-learn an optimal model initialization in the space of network weights from which the network can rapidly converge to a solution that minimizes the physics-informed loss of new tasks. This involves a task-specific fine-tuning procedure during the meta-learning stage through an inner-outer loop training framework, i.e., a bi-level optimization problem. The network weights are trained explicitly (outer loop) by optimizing the model performance over a batch of physics-informed learning tasks after one or a few gradient descent updates from the learned initialization (inner loop).
IV-D2 Neural Baldwinism
Figure 8: Schematic to illustrate new tasks arising from family of PDEs problem: S1 change to new PDE ($\gamma$ and $\delta$ parameters), source term $q$, and IC profile for $t\in[0,2]$ (same time domain as train tasks), and S2 projection to longer time domain $t\in[0,4]$. The solution for new PDE tasks (they are visually indistinguishable from the ground truth) can be quickly obtained in the order of milli-seconds by the Baldwinian-PINN [50].
Recently, a meta-learning PINN framework based on neuroevolution—dubbed Baldwinian PINN [50]—has been proposed to address the training challenges faced in conventional gradient-based meta-learning algorithms such as MAML. The Baldwinian-PINN provides a principled approach for optimizing a generic model initialization for network weights and other learning hyperparameters essential for achieving optimal performance across a distribution of physics-informed learning tasks. Interestingly, the Baldwinian-PINN limits task-specific learning to the output layer only [165, 166, 167], allowing it to solve new linear PDE problems with a 1-step pseudoinverse operation (or a few pseudoinverse steps for nonlinear PDE problems). The pseudoinverse operation is fast and exact. Unlike gradient-based MAML algorithms, the Baldwinian-PINN uses a global search strategy for neuroevolution, which can better handle the bi-level optimization problem with sparse and noisy rewards. It does not require unrolling and differentiating through the computation graph, making Baldwinian-PINN a more efficient strategy that scales well with the number of GPUs available. Fig. 7 illustrates the contrast between MAML and Baldwinian evolution for PINN [^2] Python source codes for Baldwinian-PINN are available at https://github.com/chiuph/Baldwinian-PINN..
The Baldwinian-PINN can be re-used to effectively solve new physics scenarios within a diverse task distribution with orders of magnitude improvement in speed and accuracy compared to single PINN model. The effective generalization of a Baldwinian-PINN to an entire PDE family with diverse output patterns is demonstrated in Fig. 8, on 2 task scenarios: S1 shows successful learning of solution for unseen set of PDEs, which include changes to PDE parameters, source term, and IC profile; while S2 shows effective extrapolation of solution to a longer time domain. Interestingly, the Baldwinian-PINN maintain good accuracy on tasks from S2, whereby the model first learns the PDE solution for $t\in[0,2]$, before using the learned solution $u(x,t=2)$ as new IC for solving $t\in[2,4]$. This is achievable because Baldwinian evolution does not require parameterization for the tasks, and evolved Baldwinian-PINN can generalize to new ICs, BCs, and PDE source terms. This is tricky for existing parameterized PINNs as they require interpolation across a potentially infinitely large distribution of tasks (e.g. possible BCs or ICs), further emphasizing the merits of Baldwinian evolution for physics. These encouraging results suggest that similar approaches in the spirit of Baldwinian-PINN are worthy of further consideration, such as for the continual modelling of dynamical systems through their versatility in handling ICs and ability to rapidly model and stitch time windows together with minimal error.
IV-D3 Meta-Learning Neural Architecture, Loss Function, and Optimizer
Most gradient-based meta-learning PINN approaches proposed in the literature aim to find an optimal initialization of the network weights. In the study by [168], a meta-learning framework for PINN loss function that incorporates the underlying physics of a parametrized PDE has demonstrated significant performance improvement for new tasks. Besides, a meta-learning gradient-based optimizer [169] has been recently introduced for improving hand-designed optimization rules. The meta-learned optimizer has shown significant improvement in physics-informed learning tasks compared to traditional fixed, hand-crafted optimizers, such as the Adam optimizer.
There are already a handful of studies that have proposed using EAs to find performant neural architectures and hyperparameters for a single PINN; ref. to Subsection IV-B. There is huge scope for future research and development into using EAs to jointly meta-learn model initialization, neural architecture, loss function (hyperparameters), and even the gradient-based optimizer to drastically improve convergence on a set of PINN tasks. For example, a recent work proposes the use of an EA to jointly optimize the neural architecture and network weights for a convolutional neural network-type PINN. The meta-learned PINN model integrates a physics-informed pseudoinverse computation on the final CNN layer to ensure physics compliance during prediction of a new scenario. The proposed method demonstrated enhanced generalization across diverse tasks with varying initial conditions, PDE parameters, and geometries, significantly outperforming other data-driven models like U-Nets and Fourier Neural Operators.
There exists a multitude of evolutionary computation techniques for achieving meta-learnable optimizers, loss functions, neural architectures, and model initializations over a distribution of physics-informed learning tasks. They can provide diverse training strategies for exploration to achieve effective and efficient training of a generalizable PINN. Lastly, we would like to make additional remarks on the important role of labeled data in meta-learning PINN. In general, the meta-learning of PINNs incurs a higher upfront training cost to achieve good generalization. Some proposed methods such as the Baldwinian-PINN cleverly utilize a data-centric meta-learning objective, i.e., using data MSE rather than purely physics-driven loss. Meta-learning PINN algorithms can be meaningfully guided by a few tasks with labeled data, easing optimization difficulties caused by the complex loss landscape defined by the physics-informed loss.
V Conclusion
As the influence of deep learning permeates beyond typical natural language processing or computer vision tasks, it is clear that rationalizable models that adhere to the fundamental laws of nature are a need of the hour. Instilling centuries of accumulated wisdom, in the form of mathematically expressible scientific knowledge, into otherwise naive statistical models is crucial. Research efforts in the past few years have demonstrated that PINNs are a promising approach to that end.
However, like any new technique, PINNs inevitably call for new technical innovations. Many of the approaches commonly used in training standard data-driven neural nets need to be re-evaluated, such as the method of network initialization and the choice of activation functions. The use of gradient-based optimization alone, which includes the hugely successful family of SGD algorithms, is especially worthy of reconsideration in the light of analyses summarized in this review. Identified challenges pertaining to local minima traps and vanishing gradients of physics-informed learning objectives highlight how these can prove deleterious for PINNs. These obstacles to conventional gradient-based optimization serve to outline ways in which gradient-free algorithms of computational intelligence may be uniquely advantageous as substitutes to or in synergy with SGD. Indeed, the crafting of synergistic optimization frameworks that meld gradient-based and gradient-free optimization unveils novel research questions and directions, firmly positioning scientific machine learning as an emerging area of translational interest for the computational intelligence community.
In fact, EAs have demonstrated their potential not only in PINNs but also in various approaches to solving PDEs, such as genetic programming [170, 171, 172] and reduced order models [173]. Genetic programming leverages symbolic regression to derive solutions to PDEs, offering the distinct advantage of high interpretability in the resulting solutions [174]. This capability underscores its potential for advancing the development of interpretable PDE solvers, thereby fostering enhanced understanding and usability in scientific and engineering applications. Similarly, evolution strategies have proven effective in improving the prediction accuracy of reduced order models, driven by their robust global search capabilities. The aforementioned cases highlight the significant potential of evolutionary computation in solving PDEs. We believe that this survey can serve as a valuable resource for guiding future advancements in this promising research direction.